# Asymmetric Influence of the Amplitude-Dependent Tune Shift on the Transverse Mode-Coupling Instability

### v1

@@ Abstract
At the 3 GeV ring of the MAX IV Laboratory, a fourth generation ring-based synchrotron light source, an asymmetric influence of the sign of the amplitude-dependent tune shift (ADTS) on the transverse mode-coupling instability (TMCI) has been observed. Measurements of the instability, in dedicated single-bunch experiments at low chromaticity, revealed a significant dependence of the dynamics of the instability above threshold on the sign of the ADTS. While for a negative sign of the ADTS the crossing of the instability does not lead to a loss of beam current, a positive sign results in the loss of 40% or more of the beam current at the threshold. In order to investigate the observed asymmetry systematic measurement of beam dynamics above the threshold have been conducted in combination with particle tracking simulations with mbtrack2 and theoretical calculations of the Landau damping including the ADTS. The findings indicated that this effect could become relevant in future low-emittance electron storage rings.
@@ I Introduction
The transverse mode-coupling instability (TMCI) in electron storage rings is a single-bunch transverse instability and an important collective effect which can limit the parameter space for stable operation. Especially in fourth generation light sources, this collective effect can strongly influence the achievable operation parameters. The instability depends on many beam parameters like the natural bunch length, the chromaticity and the tunes. The connection with the amplitude-dependent tune shift (ADTS) was in the past mainly investigated in the interest of mitigation of the instability by Landau damping [1, 2, 3], as the required betatron tune spread can among other sources come from the ADTS.
@@
First studies of the influence of the ADTS on the TMCI for the 3 GeV ring of the MAX IV Laboratory where presented by several of the authors in [4, 5]. While in the case of MAX IV the TMCI does not affect the standard operation, it is nevertheless important to characterize and further investigate such instabilities as with the continuous push towards more extreme operation modes and beam parameters new effects and interaction between parameters can arise and become relevant for the mitigation of such instabilities in future machines.
@@
Dedicated, systematic experiments have now been conducted in single-bunch operation showing an asymmetric influence of the sign of the ADTS on the dynamics of the vertical TMCI above threshold. While the ADTS does not seem to significantly affect the threshold current it changes the behaviour of the bunch above threshold. For values of the ADTS close to zero a partial beam loss is observed when the threshold current is crossed while slowly increasing the bunch current. For a positive sign of the ADTS this partial beam loss occurs up to higher values of the ADTS than for a negative sign, resulting in a significant asymmetry in the encountered beam loss above threshold.
@@
The beam dynamics above the TMCI threshold, can be studied at the ADTS values where not beam loss occurs. This is used to investigate the reason behind the observed asymmetry. It is observed that, as expected, the instability leads to a blow up of the bunch size and strong center-of-mass oscillations. For negative ADTS values, and at high current at high positive ADTS values, these oscillations are additionally amplitude-modulated with a much lower frequency showing a sawtooth shaped pattern.
@@
This paper will compare measurements conducted at the MAX IV Laboratory with dedicated simulations and theoretical considerations. The measurements include systematic scans of the instability threshold and the occurring current-loss as well as time-resolved measurements of the beam dynamics, in this case the center-of-mass motion and the transverse bunch size, above the threshold current.
@@ I.1 Transverse Mode-Coupling Instability
The transverse mode-coupling instability can arise when the current dependent tune shift due to the transverse impedance leads to a coupling of the coherent betatron tune with one of the neighboring head-tail mode frequencies (typically the mode -1 with a separation of $-\nu_{s}$, the synchrotron tune). The TMCI, which occurs at zero chromaticity, has, opposed to the head-tail instability, a well defined threshold current at which the growth rate increases abruptly. Figure. 1 shows the simulated mode coupling at zero chromaticity.
@@
where $\langle J\rangle$ is the average action of the particle ensemble. For our purposes, the tune shift is assumed to be proportional to the action $J$ resulting in the following definition where $b$ is the amplitude-dependent tune shift coefficient:
$\omega\left(J\right)=b\cdot J+\omega\left(0\right)\rightarrow b=\frac{\Delta\omega}{J},$ (3)
@@
Stability is then determined by evaluating the integral for solutions whose growth rate $\mathrm{Im}(V)$ is small and positive. This allows the standard exponential integral $\mathrm{Ei}$ to be used under said approximation. The dispersion relation then evaluates to
$I_{m}=-\frac{1}{b\langle J\rangle}\left[1-\zeta e^{-\zeta}(\mathrm{Ei}(\zeta)-i\pi)\right]\>\mathrm{where}\>\zeta>0$ (4)
where
@@
To include radiation damping, we could evaluate Eq. 1 for $\mathrm{Im}(V)=+1/\tau_{x}$ where $\tau_{x}$ is the radiation damping time, although the integral would then have to be evaluated numerically, slowing down the calculation significantly. In practice, the radiation damping does not make a large difference in the case of the TMCI so it is neglected.
@@ II Experimental Setup
The measurements presented in this paper were performed with the machine parameters given in Table 1 if not stated otherwise. Only a single bunch was stored in the machine to be able to use diagnostics that are not bunch-resolved such as beam position monitors (BPM) and the synchrotron light monitor (SLM). To ensure that only a pure single bunch is filled, great care has been taken to clean out residual charge from the other buckets. The total acceleration voltage in the main cavities was set to a fixed value for better comparability of different measurements. Due to the usage of a single bunch and therefore a low absolute beam current, the passive Landau cavities are not elongating the bunch. Likewise, all insertion device gaps were opened for the sake of comparability. The reference orbit for the orbit correction was set to zero-orbit without any beam line bumps laid in, to use an optic close to the design optics used in the simulations. During the measurements itself the orbit correction was only in use while changing to new settings, e.g. changing chromaticity or ADTS and was afterwards switched off. When changing settings, it was also checked that the tunes were at the standard working point. The vertical chromaticity was reduced via sextupole magnets in the vertical plane while the horizontal chromaticity remained at the value for standard operation. For adjustments of the ADTS the three octupole magnet families can be used [7].
Figure 2: BPM signal showing the amplitude modulation caused by the TMCI on the vertical center-of-mass position as function of turns.
@@ II.1 ADTS - measurement and control
The center-of-mass (COM) motion is measured via the turn-by-turn data from the beam position monitors (BPM). The position at each turn can be written out for $2^{16}$ consecutive turns. As the measurements where taken in single-bunch operation, this gives the COM position of the bunch at each turn at every BPM. Additionally, the data observed by the bunch-by-bunch system is saved with a turn-by-turn resolution as well. The coherent tunes can therefore be calculated from the Fourier transform of this data.
The amplitude-dependent tune shift (ADTS) was measured by kicking the bunch in each transverse plane individually with increasing amplitude while detecting the center-of-mass movement on the turn-by-turn BPM data for both transverse planes. The tune was calculated via the Numerical Analysis of Fundamental Frequencies (NAFF) algorithm [8, 9] based on the first 100 turns after each kick. The resulting tunes for the different kick amplitudes show the tune shift as a function of the center-of-mass displacement $\hat{x}$ at each BPM position. For the conversion between the measured maximal amplitude $\hat{x}$ at each BPM and the action $J$, it is assumed that at maximum displacement $\hat{x}$ the action $J$ can be calculated via the corresponding value of the beta-function $\beta_{\textrm{s}}$ at the position of measurement $s$, in this case the position of each BPM:
@@
Figure. 3 shows the near linear dependence of the tune shift on the action $J$ in an example measurement.
2
@@
II.2 bunch size
The bunch size is measured at the two diagnostic beam lines [10] via synchrotron light monitors (SLM) with interferometric source point imaging. Synchrotron radiation in the visible wavelength range is detected with CMOS cameras after passing though a double-slit for the horizontal plane and a diffraction obstacle in the vertical plane. The beam sizes can be calculated from the interferometric visibility in the resulting interference pattern. During the instability the bunch size is blown up to such a degree that the interference pattern is not visible anymore (Fig. 4b) and the distribution is fitted by a Gaussian. The required exposure time of the cameras to gather enough intensity, does not allow for turn-by-turn detection. Nevertheless, the exposure time is short enough ($\approx$ 1  ms) to be able to resolve the time structure of the characteristic amplitude modulation observed on the center-of-mass position (see Fig. 2) caused by the dynamics above the instability threshold. In this case, it has to be taken into account that during the exposure time the camera integrates over the observed center-of-mass oscillations providing a superposition of the center-of-mass motion and the interference pattern containing the beam size information. Furthermore, it has do be asserted that no residual bunches are present during these measurements. As visible in Fig. 4, even a small amount of charge in additional residual bunches around the main single bunch, in this case approximately 5% of the charge, can significantly influence the observed spot profile on the SLMs. As these low-charge residual bunches are below the instability threshold and therefore stable, their light shows the classical interference pattern. The intensity of the focused pattern is therefore overshadowing the smeared out spot profile from the unstable main bunch.
@@
The readout of the cameras can be triggered so that synchronised images can be taken with respect to the turn-by-turn center-of-mass motion measured with the BPMs. The synchronization was aligned with triggered kicks to the beam which can be observed in both systems (BPMs and SLMs). The timing between the camera acquisition and BPM acquisition is chosen such that the camera’s exposure time window lies roughly at three-quarters of the BPM measurement window of $2^{16}$ turns ($\approx 115$ ms). By this, the center-of-mass movement is known for some time before and after the bunch size measurement. The alignment accuracy depends on the camera exposure time used and is in the presented measurements better than $1$ ms.
@@ III Simulation tool
To simulate the beam dynamics observed, especially above the threshold current, particle tracking with the mbtrack2 [11] python code was performed.
A broadband resonator was used for the vertical impedance with a shunt impedance of 200 kOhm/m at the resonant frequency of 11.5 GHz and a quality $Q$ of 1 [12]. The mbtrack2 simulations also included a longitudinal impedance (732 Ohm at 6 GHz with $Q=1$ [13]) to account for bunch lengthening with increasing bunch current. mbtrack2 allows for the optics parameter to be read-in from an AT lattice file using pyAT. The RF voltage was set to the same value as in the measurements (see Table 1). The tune shift contribution by the ADTS is calculated in mbtrack2 based on the action $J$ (see Sec. I.1 Eq. 3). Intra-Beam Scattering (IBS) is not yet implemented in mbtrack2. The shown simulations were conducted with 50000 macro-particles and were run on the COSMOS cluster of LUNARC at Lund University.
@@
In the following the measurement and simulation results will be presented side by side and grouped by the different beam properties affected by the instability. The measurements were conducted in the vertical plane. Besides the threshold current, the beam loss at threshold, the bunch position and size as well as the betatron tune shift with current below and above the instability threshold are discussed. Additionally, theoretical calculations on the Landau damping in combination with the transverse mode-coupling instability will be discussed in context of the observed asymmetry with respect to the sign of the amplitude-dependent tune shift.
@@ IV.1 Instability Threshold
During the experimental investigations of the TMCI it was observed, that the threshold current changes depending on the beam conditions while reaching the threshold. For example, the observed threshold current was lower when the beam current slowly decayed while the beam was unstable, compared to the threshold current observed when charge was injected into a stable beam. Additionally, within a certain bunch current range, it was possible to stabilize an unstable beam with the bunch-by-bunch feedback system and the beam remained stable after switching off the feedback. Furthermore, the instability could be triggered by excitations or kicks to the beam even below the injection threshold, but not below the decaying threshold. In summary, a hysteresis effect was observed for the experimental TMCI threshold current, where a stable beam shows a higher threshold current than an already unstable or excited beam.
A possible explanation for the observed hysteresis in threshold is found when considering the effects of Intra-Beam Scattering (IBS). For beams with small transverse emittances, IBS can lead, amongst other things, to an increase in energy spread. This can be mitigated by increasing the vertical emittance either via coupling or with vertical excitations of the beam. With respect to the TMCI, IBS would have the following effect. For a stable beam the vertical emittance is small and the energy spread is increased by IBS. An increased energy spread results in an increase of the theoretical TMCI threshold, as the current-dependent tune shift is inversely proportional to the bunch length which again is proportional to the energy spread [14]. As soon as the beam becomes unstable, either by crossing the (higher) threshold or by excitation, the vertical emittance increases and the effect of the IBS is reduced leading to a reduction in energy spread. The lower energy spread finally results in a lower TMCI threshold current. This results in a hysteresis of the instability threshold depending on whether the threshold is measured starting with a stable or an unstable beam.
@@
For the following studies the threshold during injection was selected as it can be quickly measured reliably and accurately compared to the other thresholds. Furthermore, the disturbance to the stored beam caused by the used Multipole Injection Kicker (MIK) is known to be very small [15], so the observed threshold during injection should be very close to the theoretical threshold (including IBS) if the charge in a stable beam is slowly increased.
The left hand side of Fig. 6 shows the thresholds measured during injection for different ADTS coefficients. These thresholds where determined by injecting (using the MIK) into a single bunch and observing the center-of-mass movement on the BPMs. As soon as the center-of-mass movement movement grew unstable the injection was stopped and all charge in residual bunches from a non-perfect single bunch injection was cleaned. The resulting threshold currents differ as expected depending on the chromaticity. To separate the TMCI from the head-tail instability [16], the measurements where conducted either at a vertical chromaticity of zero or nearly zero chromaticity (0.05) in contrast to a chromaticity of $\approx 1.1$ during standard operation. The thresholds for both chromaticity values are very similar and lie around 2.8 mA. Additionally, measurements were conducted at a slightly negative vertical chromaticity of $-0.15$. As expected during operation with a positive momentum compaction factor and a negative chromaticity (e.g. [14]), they show a much lower threshold current of around 1 mA. The same is visible in the simulated thresholds shown on the right hand side of Fig. 6. The simulated thresholds for a chromaticity of zero and 0.05 lie both at around 3.45 mA and are higher than seen in the measurements by about 0.5 mA. At the same time, the simulated threshold for the slightly negative chromaticity matches the measurements at around 1 mA.
@@
IV.2 Beam Losses at Thresholds

### v2

@@ Abstract
At the 3 GeV ring of the MAX IV Laboratory, a fourth generation ring-based synchrotron light source, an asymmetric influence of the sign of the amplitude-dependent tune shift (ADTS) on the transverse mode-coupling instability (TMCI) has been observed. Measurements of the instability, in dedicated single-bunch experiments at low chromaticity, revealed a significant dependence of the dynamics of the instability above threshold on the sign of the ADTS. While for a negative sign of the ADTS the crossing of the instability does not lead to a loss of beam current, a positive sign results in the loss of 40% or more of the beam current at the threshold. In order to investigate the observed asymmetry, systematic measurement of beam dynamics above the threshold have been conducted in combination with particle tracking simulations with mbtrack2 and theoretical calculations of the Landau damping due to the ADTS. The findings point towards an influence of the Landau damping in combination with the low synchrotron frequency, which indicates that this effect could become relevant in future low-emittance electron storage rings.
@@ I Introduction
The transverse mode-coupling instability (TMCI) in electron storage rings is a single-bunch transverse instability and an important collective effect which can limit the parameter space for stable operation. Especially in fourth generation light sources, this collective effect can strongly influence the achievable operation parameters. The instability depends on many beam parameters like the natural bunch length, the chromaticity and the tunes. The connection with the amplitude-dependent tune shift (ADTS) was in the past investigated in the interest of mitigation of the instability by Landau damping [1, 2, 3] at proton storage rings, as the required betatron tune spread can among other sources come from the ADTS.
@@
First studies of the influence of the ADTS on the TMCI for the 3 GeV ring of the MAX IV Laboratory where presented by several of the authors in [4, 5]. While in the case of MAX IV the TMCI does not affect routine operation, it is nevertheless important to characterize and further investigate such instabilities as with the continuous push towards more extreme operation modes and beam parameters new effects and interaction between parameters can arise and become relevant for the mitigation of such instabilities in future machines.
@@
Dedicated, systematic experiments have now been conducted in single-bunch operation showing an asymmetric influence of the sign of the ADTS on the dynamics of the vertical TMCI above threshold. While the ADTS does not seem to significantly affect the threshold current, it changes the behaviour of the bunch above threshold. For values of the ADTS close to zero, a partial beam loss is observed when the threshold current is crossed while slowly increasing the bunch current. For ADTS with a large absolute value, on the other hand, beam loss is not observed as the threshold is crossed and the instability leads solely to oscillations of the bunch centre-of-mass and a beam size blow-up. For a positive sign of the ADTS, this partial beam loss occurs up to higher values of the ADTS than for a negative sign, resulting in a significant asymmetry in the encountered beam loss above threshold.
@@
At ADTS values where no beam loss occurs, the beam dynamics above the TMCI threshold can be studied. This is used to investigate the observed asymmetry further. It is observed that, as expected, the instability leads to a blow-up of the bunch size and strong center-of-mass oscillations. For negative ADTS values, and at high current at high positive ADTS values, these oscillations are additionally amplitude-modulated with a much lower frequency showing a sawtooth shaped pattern.
@@
This paper compares measurements conducted at the MAX IV Laboratory with dedicated simulations and theoretical considerations. The measurements include systematic scans of the instability threshold and the occurring current loss as well as time-resolved measurements of the beam dynamics, in this case the center-of-mass motion and the transverse bunch size, above the threshold current. The simulations consist of particle tracking with the mbtrack2 code [6] which allows the inclusion of the amplitude-dependent tune shift. Theoretical calculation considering transverse mode-coupling and Landau damping due to the ADTS were conducted as described in the next section and section IV.4.
@@ I.1 Transverse Mode-Coupling Instability
The transverse mode-coupling instability can arise when the current dependent tune shift due to the transverse impedance leads to a coupling of the coherent betatron tune with one of the neighboring head-tail mode frequencies (typically the mode -1 with a separation of $-\nu_{s}$, the synchrotron tune). The TMCI, which occurs at zero chromaticity, has, opposed to the head-tail instability, a well defined threshold current at which the growth rate increases abruptly. Figure 1 shows the simulated mode coupling at zero chromaticity.
@@
where $\langle J\rangle$ is the average action of the particle ensemble. For our purposes, the tune shift is assumed to be proportional to the action $J$ resulting in the following definition where $b$ is the amplitude-dependent
$\Delta\nu\left(J\right)=b\cdot J$ (3)
@@
Evaluating the integral in Eq. 1 gives
$I_{m}=-\frac{1}{b\langle J\rangle}\left[1+\zeta e^{-\zeta}E_{1}(-\zeta)\right]\;\text{where}\;\zeta>0$ (4)
and
@@
Solutions with small and positive growth rates $\mathrm{Im}(V)>0$ lie on the boundary of stability. To include radiation damping, we could set $\mathrm{Im}(V)=+1/\tau_{x}$ where $\tau_{x}$ is the radiation damping time. Here, however, we neglect radiation damping as it does not make a large difference in the case of the TMCI. An added benefit is that Eq. 4 is then scalable by $b$ and so only needs to be evaluated twice, once for $b>0$ and once for $b<0$.
@@ II Experimental Setup
Horizontal beta function / m 0.8 to 9.8
The measurements presented in this paper were performed with the machine parameters given in Table 1 if not stated otherwise. Only a single bunch was stored in the machine to be able to use diagnostics that are not bunch-resolved such as beam position monitors (BPM) and the synchrotron light monitor (SLM). To ensure that only a pure single bunch is filled, great care has been taken to clean out residual charge from the other buckets. The used BPMs and the used SLM sit at different position in the ring, and therefore at different values of the beta function. To make the bunch sizes and center-of-mass positions shown within this publication comparable, all measurements are scaled to correspond to a measurement at a beta function of 16.0 m. This value was chosen as it is the beta function at the position of the used SLM and is very close to the maximal vertical beta function of 16.1 m. To translate the measured values to this position, they are scaled with the ratio of the square roots of the goal beta function value and the original value at the place of measurement.
The total acceleration voltage in the main cavities was set to a fixed value for better comparability of different measurements. Due to the usage of a single bunch and therefore a low absolute beam current, the passive Landau cavities are not elongating the bunch. Likewise, all insertion device gaps were opened for the sake of comparability. The reference orbit for the orbit correction was set to zero-orbit without any beamline bumps [^1] Beamline bumps refer to deliberate deviations in the closed orbit at the position of IDs to help synchrotron radiation alignment for beamlines. laid in, to use an optic close to the design optics used in the simulations. During the measurements, the orbit correction was only in use while changing to new settings, e.g. changing chromaticity or ADTS and was afterwards switched off. When changing settings, it was also checked that the tunes were at the routine working point. The vertical chromaticity was reduced via sextupole magnets in the vertical plane while the horizontal chromaticity remained at the value for routine operation of around 1.1. For adjustments of the ADTS three octupole magnet families are available [8].
Figure 2: BPM signal showing the amplitude modulation caused by the TMCI on the vertical center-of-mass position as function of turns for an ADTS coefficient of $b=-6868/\mathrm{m}$ and 4.5 mA.
@@ II.1 ADTS - Measurement and Control
The center-of-mass (COM) motion is measured via the turn-by-turn data from the BPMs. The position at each turn can be written out for $2^{16}$ consecutive turns. As the measurements where taken in single-bunch operation, this gives the COM position of the bunch at each turn at every BPM. The coherent tunes can therefore be calculated from the Fourier transform of this data.
The amplitude-dependent tune shift was measured by kicking the bunch in each transverse plane individually with increasing amplitude while detecting the center-of-mass movement on the turn-by-turn BPM data for both transverse planes. The tune was calculated via the Numerical Analysis of Fundamental Frequencies (NAFF) algorithm [9, 10] based on the first 100 turns after each kick. The resulting tunes for the different kick amplitudes show the tune shift as a function of the center-of-mass displacement $\hat{x}$ at each BPM position. For the conversion between the measured maximal amplitude $\hat{x}$ at each BPM and the action $J$, it is assumed that at maximum displacement $\hat{x}$ the action $J$ can be calculated via the corresponding value of the beta-function $\beta_{\textrm{s}}$ at the position of measurement $s$, in this case the position of each BPM:
@@
Figure 3 shows the near-linear dependence of the tune shift on the action $J$ in an example measurement.
@@
II.2 Transverse Bunch Size
The transverse bunch size is measured at the two diagnostic beam lines [11] via synchrotron light monitors (SLM) with interferometric source point imaging. Synchrotron radiation in the visible wavelength range is detected with CMOS cameras after passing though a double-slit for the horizontal plane and a diffraction obstacle in the vertical plane. The beam sizes can be calculated from the interferometric visibility in the resulting interference pattern [11]. During the instability, the bunch size is blown up to such a degree that the interference pattern is not visible anymore (Fig. 4b) and the distribution is fitted by a Gaussian. The required exposure time of the cameras to gather enough intensity, does not allow for turn-by-turn detection. Nevertheless, the exposure time is short enough ($\approx$ 1  ms) to allow resolving the time structure of the characteristic amplitude modulation observed on the center-of-mass position (see Fig. 2) caused by the dynamics above the instability threshold. In this case, it has to be taken into account that during the exposure time the camera integrates over the observed center-of-mass oscillations providing a superposition of the center-of-mass motion and the interference pattern containing the beam size information. Furthermore, it has do be asserted that no residual bunches are present during these measurements. As visible in Fig. 4, even a small amount of charge in additional residual bunches around the main single bunch, in this case approximately 5% of the charge, can significantly influence the observed spot profile on the SLMs. As these low-charge residual bunches are below the instability threshold and therefore stable, their light shows the classical interference pattern. The intensity of the focused pattern is therefore overshadowing the smeared-out spot profile from the unstable main bunch.
@@ II.3 Synchronised Measurements
The acquisition by the cameras can be triggered so that synchronised images can be taken relative to the turn-by-turn center-of-mass motion measured with the BPMs. The synchronization was aligned with triggered kicks from a pulsed magnet to the beam which can be observed in both systems (BPMs and SLMs). The timing between the camera acquisition and BPM acquisition is chosen such that the camera’s exposure time window lies roughly at three-quarters of the BPM measurement window of $2^{16}$ turns ($\approx 115$ ms). By this, the center-of-mass movement is known for some time before and after the bunch size measurement. The alignment accuracy depends on the camera exposure time used and is in the presented measurements better than $1$ ms.
@@ III Simulation tool
To simulate the beam dynamics observed, especially above the threshold current, we performed particle tracking with the mbtrack2 [6] python code.
A broadband resonator was used for the vertical impedance with a shunt impedance of 200 kOhm/m at the resonant frequency of 11.5 GHz and a quality $Q$ of 1 [12]. The mbtrack2 simulations also included a longitudinal impedance (732 Ohm at 6 GHz with $Q=1$ [13]) to account for bunch lengthening with increasing bunch current. mbtrack2 allows for the optics parameter to be read-in from an AT lattice file using pyAT. The RF voltage was set to the same value as in the measurements (see Table 1) and synchrotron radiation effects were included in the simulations. The tune shift contribution by the ADTS is calculated in mbtrack2 based on the action $J$ (see Sec. I.1 Eq. 3). Intra-Beam Scattering (IBS) is not yet implemented in mbtrack2. The simulations were conducted with 50000 macro-particles and were run on the COSMOS cluster of LUNARC at Lund University. The center-of-mass amplitudes and bunch sizes are calculated in mbtrack2 once per turn and for a value of the beta function of 7.23 m corresponding approximately to the average of the beta function along the ring. Therefore, all simulated COM amplitudes and bunch sizes were scaled to the same reference as the measurements, a beta function of 16.0 m.
@@
In the following the measurement and simulation results are presented side by side and grouped by the different beam properties affected by the instability. The measurements were conducted in the vertical plane. Besides the threshold current, the beam loss at threshold, the bunch position and size as well as the betatron tune shift with current below and above the instability threshold are discussed. Additionally, theoretical calculations on the Landau damping in combination with the transverse mode-coupling instability will be discussed in the context of the observed asymmetry with respect to the sign of the amplitude-dependent tune shift.
@@ IV.1 Instability Threshold
We observed that the TMCI threshold current changes depending on the beam conditions while reaching the threshold. For example, the threshold current was lower when the beam current slowly decayed while the beam was unstable, compared to the threshold current when charge was injected into a stable beam. Additionally, within a certain bunch current range, it was possible to stabilize an unstable beam with the bunch-by-bunch feedback system and the beam remained stable after switching off the feedback. Furthermore, the instability could be triggered by excitations or kicks to the beam even below the injection threshold, but not below the decaying threshold. In summary, we observed an hysteresis effect on the TMCI threshold current, where a stable beam shows a higher threshold current than an already unstable or excited beam.
A possible explanation for the threshold hysteresis is Intra-Beam Scattering. For beams with small transverse emittances, IBS can lead, amongst other things, to an increase in energy spread. This can be mitigated by increasing the vertical emittance either via coupling or with vertical excitations of the beam. With respect to the TMCI, IBS would have the following effect. For a stable beam the vertical emittance is small and the energy spread is increased by IBS. An increased energy spread results in an increase of the theoretical TMCI threshold, as the current-dependent tune shift is inversely proportional to the bunch length which again is proportional to the energy spread [14]. As soon as the beam becomes unstable, either by crossing the (higher) threshold or by excitation, the vertical emittance increases and the effect of the IBS is reduced leading to a reduction in energy spread. The lower energy spread finally results in a lower TMCI threshold current. This results in a hysteresis of the instability threshold depending on whether the threshold is measured starting with a stable or an unstable beam.
@@
For the following studies, the threshold during injection was selected as it can be quickly measured reliably and accurately compared to the other thresholds. Furthermore, the disturbance to the stored beam caused by the Multipole Injection Kicker (MIK) is known to be very small [15], so the observed threshold during injection should be very close to the theoretical threshold (including IBS) if the charge in a stable beam is slowly increased.
The left hand side of Fig. 6 shows the thresholds measured during injection for different ADTS coefficients. These thresholds where determined by injecting (using the MIK) into a single bunch and observing the center-of-mass movement on the BPMs. As soon as the center-of-mass movement grew unstable the injection was stopped and all charge in residual bunches from a non-perfect single bunch injection was cleaned. The resulting threshold currents differ as expected depending on the chromaticity. To separate the TMCI from the head-tail instability [16], the measurements where conducted either at a vertical chromaticity of zero or nearly zero chromaticity (0.05) in contrast to a chromaticity of $\approx 1.1$ during routine operation. The thresholds for both chromaticity values are very similar and lie around 2.8 mA. Additionally, measurements were conducted at a slightly negative vertical chromaticity of $-0.15$. As expected during operation with a positive momentum compaction factor and a negative chromaticity (e.g. [14]), they show a much lower threshold current of around 1 mA. The same is visible in the simulated thresholds shown on the right hand side of Fig. 6. The simulated thresholds for a chromaticity of zero and 0.05 lie both at around 3.45 mA and are higher than seen in the measurements by about 0.5 mA. At the same time, the simulated threshold for the slightly negative chromaticity matches the measurements at around 1 mA.
@@
IV.2 Beam Losses at Threshold